# Score-based Generative Priors Guided Model-driven Network for MRI Reconstruction

### v1

@@ Abstract
Score matching with Langevin dynamics (SMLD) method has been successfully applied to accelerated MRI. However, the hyperparameters in the sampling process require subtle tuning, otherwise the results can be severely corrupted by hallucination artifacts, particularly with out-of-distribution test data. In this study, we propose a novel workflow in which SMLD results are regarded as additional priors to guide model-driven network training. First, we adopted a pretrained score network to obtain samples as preliminary guidance images (PGI) without the need for network retraining, parameter tuning and in-distribution test data. Although PGIs are corrupted by hallucination artifacts, we believe that they can provide extra information through effective denoising steps to facilitate reconstruction. Therefore, we designed a denoising module (DM) in the second step to improve the quality of PGIs. The features are extracted from the components of Langevin dynamics and the same score network with fine-tuning; hence, we can directly learn the artifact patterns. Third, we designed a model-driven network whose training is guided by denoised PGIs (DGIs). DGIs are densely connected with intermediate reconstructions in each cascade to enrich the features and are periodically updated to provide more accurate guidance. Our experiments on different sequences revealed that despite the low average quality of PGIs, the proposed workflow can effectively extract valuable information to guide the network training, even with severely reduced training data and sampling steps. Our method outperforms other cutting-edge techniques by effectively mitigating hallucination artifacts, yielding robust and high-quality reconstruction results.
Index Terms:
@@
MRI reconstruction, model-driven, diffusion model, compressed sensing, denoising score matching.
I Introduction
@@
Magnetic Resonance Imaging (MRI) inherently requires long scanning times, leading to patient discomfort and degradation of image quality. Undersampling the k-space data can proportionally reduce the scanning time but results in aliasing artifacts in the reconstructed images. The conventional methods for reconstructing high-quality images from a small subset of k-space data are mainly parallel imaging (PI) [1, 2] and compressed sensing (CS) [3, 4, 5]. In recent years, deep-learning (DL)-based methods have been successfully introduced in the field of MRI reconstruction and have shown convincing improvements over conventional methods [6, 7]. For DL-based reconstruction, the proposed reconstruction models can be categorized into data-driven and model-driven unrolled networks. Part of the data-driven model was designed in an end-to-end manner to fit the mapping between the input measurements and reconstruction targets [8, 9, 10]. However, these methods still present problems such as lack of interpretability, and a relatively large dataset is required to fit the mapping. Recently, some data-driven networks have focused on distribution-learning-based reconstruction processes using diffusion models [11, 12, 13, 14, 15, 16, 17]. Among them, denoising diffusion probabilistic models (DDPMs) [18, 19] and score matching [20, 21] with Langevin dynamics (SMLD) methods, which correspond to Varience Preserving (VP) Stochastic differential equations (SDEs) and Variaence Exploding (VE) SDEs, respectively [22], have shown great potential in solving inverse problems. They have also been successfully extended to the MRI reconstruction field [23, 24, 22, 25]. However, the training (i.e., forward) and sampling (i.e., reverse) processes require careful tuning, otherwise the sampling results can be corrupted by hallucination artifacts, particularly with out-of-distribution test data.
@@
Figure 1: Distribution-shift reconstruction results of a pretrained score network on a knee MRI dataset. The score network was trained on a brain MRI dataset. Images in odd columns are objective reality and those in even columns are sampling results.
@@
Model-driven unrolled networks are built in a different manner: conventional optimization algorithms are unrolled and realized in deep networks, and stacked cascades are used to mimic the iterative steps in the optimization procedure. Compared with data-driven methods, unrolled networks can be better interpreted as those in which conventional priors can be appropriately incorporated, and they have shown convincing results in recent years. Although stacking more cascades can improve the reconstruction performance, large-scale backbone networks can hardly be applied to unrolled networks, considering the limitations of GPU memory, and a trade-off must be established. However, data- and model-driven methods attempt to recover the same target images along different paths, and the complementarity between these methods should be considered. Moreover, we observed the convincing potential of SMLD methods for MRI reconstruction. Current score-matching networks require retraining when applied to different datasets and we observe artifacts on the sampled images without network retraining and subtle tuning of hyperparameters, as shown in Fig. 1. However, we believe that the artifact-affected images can still provide information additional to the known undersampled measurements if the artifacts can be appropriately eliminated, thereby compensating for missing measurements.
@@
Instead of focusing on optimizing the forward and reverse process in SMLD methods, we aim to design a model-driven network which can extract valuable information from naive sampling process and achieve high-quality results robustly. We proposed a novel DL-based hybrid MRI reconstruction workflow, consisting of the following three main steps: (1) Sampling: Using known undersampled raw k-space data, we first obtain samples using SMLD methods. We bypassed all the tuning steps in SMLD methods, such as tuning the hyperparameters or retraining the score networks on our dataset. Instead, we adopted a pretrained score network (PSN) and used the naive samples as preliminary guidance images (PGI) for the follow-up steps. A PSN can be trained using out-of-distribution data. (2) Denoising: We believe that these preliminary reconstructions can provide additional information to guide the reconstruction process through appropriate denoising, leading to a better performance. Considering that the PGIs have relatively low quality and that we observe similar aliasing artifacts in the images, we designed a denoising module (DM) to remove the artifacts from the PGIs. We extracted features from the same score network after fine-tuning and concatenated them with the inputs in Langevin dynamics, which directly produces artifacts, to learn the noise pattern. Then, we used denoised PGIs (DGIs) as higher-quality guidelines for further reconstruction. (3) Guided reconstruction: We designed a densely connected unrolled network to reconstruct the MR images from the undersampled measurements. DGIs are simultaneously input with measurements and periodically updated in the network to guide network training. The DM and unrolled cascades were trained end-to-end.
@@
The experimental results show that although the quality of PGIs is even lower than that of the undersampled measurements, the artifacts were effectively eliminated by the DMs. Meanwhile, the score network can not only help in the sampling process but also cooperate with the unrolled network and improve the denoising performance of the DMs. The proposed reconstruction workflow outperformed state-of-the-art methods through the guidance of DGIs.
@@
    We propose a novel DL-based MRI reconstruction workflow, and low-quality samples obtained using SMLD method are used as guidance during reconstruction.
  2. 2.
    We design an unrolled network, and the initially obtained samples are denoised and periodically updated to guide network training.
  3. 3.
@@
    Experimental results show that without tuning the sampling steps or retraining the score network on target dataset, the low-quality samples can guide the proposed network to obtain robust and high-quality reconstructions, which outperforms the cutting-edge methods.
II Related Works
II-A Conventional approaches for accelerated MRI
@@
For the PI, multiple receiver coils are equipped to obtain raw k-space data simultaneously, and the missing data are recovered by exploiting redundancies among the data from different coils. Sensitivity encoding (SENSE)-type [26] methods work in the image domain, whereas others directly interpolate missing data points in k-space, such as simultaneous acquisition of spatial harmonics (SMASH) [27] and generalized autocalibration partial parallel acquisition (GRAPPA) [28]. However, the reconstructions in PI approaches deteriorate at high acceleration factors. In CS, sparsity-based priors are routinely incorporated to narrow the solution space in the objective functions, and iterative optimization algorithms are designed to approach artifact-free images [29, 30, 31, 32, 33, 34]. However, conventional methods depend heavily on handcrafted procedures, and the iterative steps are time-consuming.
Figure 2: (a) The workflow of the proposed SGM-Net. (b) The network structure of the sampling and denoising steps. (c) The network structure of the $k$th cascade.
II-B Data-driven networks for MRI reconstruction
@@
For data-driven networks trained end-to-end, the mapping between the undersampled measurements and reconstructed images was established. For example, early studies adopted deep convolutional neural networks (CNNs) to reconstruct MR images from zero-filled inputs [35, 36]. The authors in [37] trained the magnitude and phase networks separately in a deep residual network to remove aliasing artifacts. In [38], a complex-valued deep residual network with a higher performance and smaller size was proposed. In addition, GAN-[39, 40]-based networks have shown convincing results [41, 42, 43]. In addition to realizing image domain end-to-end learning, some studies have proposed direct learning mapping in k-space [44] or exploiting the complementarity between the image and k-space domains [9].
@@
Distribution learning is another category of data-driven networks. For example, in [45] the authors used pre-trained generative models to present data distributions. In [23], the authors used score matching to estimate the data distribution, and samples were then generated using annealed Langevin dynamics. Subsequently, the authors proposed improved skills in [24] for higher stability and better performance, and exact log-likelihood computation was supported in [22]. [25] first successfully applied CS-based generative models to overcome the MRI reconstruction problem and showed convincing results in out-of-distribution samples. In [16], the authors further optimized the sampling stage by automatically tuning the hyperparameters using Stein’s unbiased risk estimator (SURE). The authors in [17] proposed a high-frenquency-based diffusion process and accelerated the reverse process. A wavelet-improved technique was proposed in [15] to stably train the score network. In the reverse process, the authors further designed regularization constraints to enhance the robustness. Some works extended the idea of cold diffusion [46] to MRI reconstruction, and the process of adding Gaussian noise is replaced with undersampling operation [47]. However, network retraining is necessary when changing the acquisition parameters, and large datasets are required to fit the mapping [48].
II-C Unrolled networks for MRI reconstruction
@@
Unrolled networks, also known as model-driven networks, are constructed by unrolling iterative optimization steps into deep networks [49, 50, 51, 52, 53]. For example, algorithms such as the alternating direction method of multipliers (ADMM) [49], iterative shrinkage-thresholding algorithm (ISTA) [50], variable splitting (VS) [52], and approximate message passing (AMP) have been successfully applied to solve the objective functions in MRI reconstruction. In [53], the authors extended a previous variational network [51] to learn end-to-end, and sensitivity maps (SM) were estimated in the network. To further improve the accuracy of the SM, different studies [54, 55] have explored ways to simultaneously update the SMs with the reconstructed images. In [56], convolutional blocks were used to extract high-resolution features and transformer blocks have been introduced to refine low-resolution features. The authors of [57] used pre-acquired intra-subject MRI modalities to guide the reconstruction process. In [58], the authors designed a transformer-enhanced network, incorporating a regularization term on the error maps within the residual image domain. Some works solved the optimization problem with structured low-rank algorithm [59, 60, 61]. The authors in [62] designed a deep low-rank and sparse network which explored 1D convolution during training. However, the complementarity between unrolled networks and data-driven methods has not been investigated; in this study, we propose a novel concept in which we treat naive sampling results from SMLD methods as guidance and unrolled cascades as denoising and refining processes.
III Background
@@
$\mathcal{E}(x)=(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...),i=1,2,...,C$ (3)
@@
$\mathcal{R}(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...)=\sum_{i=1}^{C}\mathcal{S}_{i}x$ (4)
@@
It is ill-posed to recover missing data directly from $y$. Hence, a regularization term $\lambda\mathcal{U}(x)$ is routinely added to narrow the solution space in the CS, which is
$\hat{x}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x-y\|_{2}^{2}+\lambda\mathcal{U}(x)$ (6)
For example, $\lambda\Psi(x)$ is a common regularization term that imposes the sparsity constraint. Then, gradient descent can be used to solve the problem iteratively as follows:
@@
where $\alpha^{t}$ is the step size in the $t$th iteration and $\psi(x^{t})$ is the gradient of $\Psi(x)$.
IV Methods
We proposed a score-based generative priors guided model-driven Network (SGM-Net), which is illustrated in Fig. 2. The whole workflow is illustrated in Fig. 2 (a), which comprises three parts: sampling (green components), denoising (blue components), and guided reconstruction (yellow components). The preliminary samples (that is, $x_{T}$ in Fig. 2 (a)) were first obtained using Langevin dynamics and then used as PGIs for the follow-up steps. The gradients of the data distribution were obtained using PSN. Instead of carefully training and tuning the SMLD process to obtain high-quality samples, we sampled $x_{T}$ from an out-of-distribution pretrained network with fitted hyperparameters. Then, we treat the artifacts in $x_{T}$ as removable and propose a DM to eliminate noise and enhance $x_{T}$. The information available from the sampling steps was integrated and fed into the DM, as shown in Fig. 2 (b). Finally, DGIs were used to guide an unrolled network, as shown in Fig. 2 (c). Details of the proposed method are presented in the following subsections.
IV-A Score matching and Langevin dynamics
The first step of the proposed workflow is to obtain PGIs with SMLD to guide the subsequent steps. In this study, we adopted the PSN proposed in [23], which was successfully extended to MRI reconstruction in [25]. To improve the generality of our workflow, the score network was pretrained with out-of-distribution data and tedious tuning and retraining steps were bypassed. The network, remarked as $s_{\theta}$, is a four-cascade RefineNet [63] with dilated convolutions and instance normalization, which is used to approach the score function of $p(x)$ (that is, $\bigtriangledown_{x}\log p(x)$) and we have $s_{\theta}\approx\bigtriangledown_{x}\log p(x)$. Average pooling layers were adopted to replace all maximum pooling layers in RefineNet, and ELU was adopted as the activation function. Then, the annealed Langevin method-based sampling process in MRI can be represented as

### v2

@@ Abstract
Score matching with Langevin dynamics (SMLD) method has been successfully applied to accelerated MRI. However, the hyperparameters in the sampling process require subtle tuning, otherwise the results can be severely corrupted by hallucination artifacts, especially with out-of-distribution test data. To address the limitations, we proposed a novel workflow where naive SMLD samples serve as additional priors to guide model-driven network training. First, we adopted a pretrained score network to generate samples as preliminary guidance images (PGI), obviating the need for network retraining, parameter tuning and in-distribution test data. Although PGIs are corrupted by hallucination artifacts, we believe they can provide extra information through effective denoising steps to facilitate reconstruction. Therefore, we designed a denoising module (DM) in the second step to coarsely eliminate artifacts and noises from PGIs. The features are extracted from a score-based information extractor (SIE) and a cross-domain information extractor (CIE), which directly map to the noise patterns. Third, we designed a model-driven network guided by denoised PGIs (DGIs) to further recover fine details. DGIs are densely connected with intermediate reconstructions in each cascade to enrich the information and are periodically updated to provide more accurate guidance. Our experiments on different datasets reveal that despite the low average quality of PGIs, the proposed workflow can effectively extract valuable information to guide the network training, even with severely reduced training data and sampling steps. Our method outperforms other cutting-edge techniques by effectively mitigating hallucination artifacts, yielding robust and high-quality reconstruction results.
keywords:
@@
MRI reconstruction , model-driven , diffusion model , compressed sensing , denoising score matching
journal: Biomedical Signal Processing and Control
[…]
1 Introduction
@@
Magnetic Resonance Imaging (MRI) inherently requires long scanning time, leading to patient discomfort and degradation of image quality. Undersampling the k-space data can proportionally reduce the scanning time but results in aliasing artifacts in the reconstructed images. The conventional methods for reconstructing high-quality images from a small subset of k-space data are mainly based on parallel imaging (PI) [1, 2] and compressed sensing (CS) [3, 4, 5]. In recent years, deep-learning (DL)-based methods have been successfully introduced in the field of MRI reconstruction and have shown convincing improvements over conventional methods [6, 7]. Early networks are designed in an end-to-end manner to fit the mapping between the input measurements and reconstruction targets [8, 9]. However, these methods still present problems such as lack of interpretability, and a relatively large dataset is required to fit the mapping. Recently, some studies have focused on distribution-learning-based reconstruction processes using diffusion models [10, 11, 12, 13, 14, 15, 16]. Among them, denoising diffusion probabilistic models (DDPMs) [17, 18] and score matching [19, 20] with Langevin dynamics (SMLD) methods, which correspond to variance preserving (VP) stochastic differential equations (SDEs) and variance exploding (VE) SDEs, respectively [21], have shown great potential in solving inverse problems. They have also been successfully extended to the MRI reconstruction field [22, 23, 21, 24]. However, the training (i.e., forward) and sampling (i.e., reverse) processes require careful tuning, otherwise the sampling results can be corrupted by hallucination artifacts, particularly with out-of-distribution test data.
@@
Figure 1: Distribution-shift reconstruction results of a pretrained score network on a knee MRI dataset. The score network was trained on a brain MRI dataset. Images in odd columns are ground truth and those in even columns are sampling results.
@@
Model-driven networks are built in a different manner: conventional optimization algorithms are unrolled and realized in deep networks, and stacked cascades are used to mimic the iterative steps in the optimization procedure. Compared with the other methods, unrolled networks offer better interpretability. Enhanced reconstruction performance is observed with learning based parameters and regularizers in unrolled networks compared with traditional optimization procedure. Although stacking more cascades can improve the reconstruction performance, large-scale backbone networks can hardly be applied to unrolled networks due to the limitations of GPU memory, and a trade-off must be established. Considering that distribution-learning-based and model-driven methods attempt to recover the same target images along different paths, the complementarity between these methods should be considered. Current score-matching networks require retraining to achieve superior results across varying datasets. We observed artifacts on the sampled images without network retraining and tuning of hyperparameters, as shown in Fig. 1. Despite these artifacts, we can observe that the artifact-affected images still provide information additional to the known undersampled measurements. Therefore, if the artifacts can be appropriately eliminated, naive samples from SMLD methods can compensate for missing measurements and provide complementary information to guide model-driven network training.
@@
To this end, instead of focusing on optimizing the forward and reverse process in SMLD methods, we aimed to design a model-driven network which can extract valuable information from naive sampling process and achieve high-quality results robustly. We proposed a novel DL-based hybrid MRI reconstruction workflow, consisting of the following three main steps: (1) Sampling: Using known undersampled raw k-space data, we first obtained initial samples using SMLD methods. We bypassed all the tuning steps in SMLD methods, such as tuning the hyperparameters or retraining the score networks on our dataset. Instead, we adopted a pretrained score network (PSN) and used the naive samples as preliminary guidance images (PGI) for the follow-up steps. A PSN can be trained using out-of-distribution data. (2) Denoising: We believe that these preliminary reconstructions can provide additional information to guide the reconstruction process through appropriate denoising, leading to better performance. Considering that the PGIs have relatively low quality and we observed similar artifacts in the images, we designed a denoising module (DM) to remove the artifacts and noises from the PGIs. We designed a score-based information extractor (SIE) and a cross-domain information extractor (CIE) to build mappings from extracted features to the noise patterns. Then, we used denoised PGIs (DGIs) as higher-quality guidelines for further reconstruction. (3) Guided reconstruction: We designed a densely connected unrolled network to reconstruct the MR images from the undersampled measurements. DGIs are simultaneously input with measurements and periodically updated in the network to guide network training. The DM and unrolled cascades were trained end-to-end. Finally, the proposed workflow essentially reconstruct MR images in a coarse-to-fine manner. The denoising steps primarily restore low frequency patterns, while the guided reconstruction steps further recover fine details, achieving superior image quality.
@@
The experimental results show that although the average quality of PGIs is even lower than that of the undersampled measurements, the artifacts were effectively eliminated by the DMs. Meanwhile, the score network can not only help in the sampling process but also cooperate with the unrolled network and improve the denoising performance of the DMs. Finally, the proposed reconstruction workflow outperformed state-of-the-art methods through the guidance of DGIs.
@@
    We propose a novel DL-based MRI reconstruction workflow, and low-quality samples obtained using naive SMLD method can be leveraged as guidance during reconstruction.
  • 2.
    We propose a DM with SIE and CIE blocks which can coarsely remove artifacts and noises in the SMLD samples, and an unrolled network is proposed to further recover fine details.
  • 3.
@@
    Experimental results demonstrate that without tuning the sampling steps or retraining the score network on target dataset, the low-quality samples can guide the proposed network to obtain robust and high-quality reconstructions, which outperforms the cutting-edge methods.
2 Related Works
@@
For PI methods, multiple receiver coils are equipped to obtain raw k-space data simultaneously, and the missing data are recovered by exploiting redundancies among the data from different coils. Sensitivity encoding (SENSE)-type [25] methods work in the image domain, whereas some others directly interpolate missing data points in k-space, such as simultaneous acquisition of spatial harmonics (SMASH) [26] and generalized autocalibration partial parallel acquisition (GRAPPA) [27]. However, the reconstructions in PI approaches deteriorate at high acceleration factors. In CS, sparsity-based priors are routinely incorporated to narrow the solution space in the objective functions, and iterative optimization algorithms are designed to approach artifact-free images [28, 29, 30, 31, 32, 33]. However, conventional methods depend heavily on handcrafted procedures, and the iterative steps are time-consuming.
@@
For deep networks trained end-to-end, the mapping between the undersampled measurements and reconstructed images are established. For example, early studies adopted deep convolutional neural networks (CNNs) to reconstruct MR images from zero-filled inputs [34, 35]. The authors in [36] trained the magnitude and phase networks separately in a deep residual network to remove aliasing artifacts. In [37], a complex-valued deep residual network with a higher performance and smaller size was proposed. In addition to realizing image domain end-to-end learning, some studies have proposed direct learning mapping in k-space [38] or exploiting the complementarity between the image and k-space domains [9, 39]. In [40], the authors proposed an unsupervised framework which learn from image and frequency domain in a multi-stage manner.
@@
Some other methods direct learned an image prior in generative models such as generative adversarial networks (GAN) [41, 42, 43, 44, 45, 46]. For example, in [47] the authors used pre-trained generative models to present data distributions. More recently, diffusion models have emerged as new state-of-the-art generative models. In [22], the authors used score matching to estimate the data distribution, and samples were then generated using annealed Langevin dynamics. Subsequently, the authors proposed improved skills in [23] for higher stability and better performance, and exact log-likelihood computation was supported in [21]. [24] first successfully applied CS-based generative models to overcome the MRI reconstruction problem. In [15], the authors further optimized the sampling stage by automatically tuning the hyperparameters using Stein’s unbiased risk estimator (SURE). The authors in [16] proposed a high-frenquency-based diffusion process and accelerated the reverse process. A wavelet-improved technique was proposed in [14] to stably train the score network. In the reverse process, the authors further designed regularization constraints to enhance the robustness. Some works extended the idea of cold diffusion [48] to MRI reconstruction, and the process of adding Gaussian noise is replaced with undersampling operation [49]. However, these networks generally required large training datasets [50] and suffer from the lack of interpretability. Moreover, tuning the training and sampling process of diffusion models is non-trivial and crucial for achieving robust reconstruction outcomes.
@@
Unrolled networks, also known as model-driven networks, are constructed by unrolling iterative optimization steps into deep networks [51, 52, 53, 54, 55, 56, 57]. For example, algorithms such as the alternating direction method of multipliers (ADMM) [51], iterative shrinkage-thresholding algorithm (ISTA) [52], variable splitting (VS) [54], and approximate message passing (AMP) have been successfully applied to solve the objective functions in MRI reconstruction. In [55], the authors extended a previous variational network [53] to learn end-to-end, and sensitivity maps (SM) were estimated in the network. To further improve the accuracy of the SM, different studies [58, 59] have explored ways to simultaneously update the SMs with the reconstructed images. In [60], convolutional blocks were used to extract high-resolution features and transformer blocks have been introduced to refine low-resolution features. The authors of [61] used pre-acquired intra-subject MRI modalities to guide the reconstruction process. In [62], the authors designed a transformer-enhanced network, incorporating a regularization term on the error maps within the residual image domain. Some works solved the optimization problem with structured low-rank algorithm [63, 64, 65]. The authors in [66] designed a deep low-rank and sparse network which explored 1D convolution during training. In [67], ADMM was utilized to unroll a half-quadratic variable splitting based optimization process, and a dilated-convolution model was designed to predict the Lagrange multipliers. However, the complementarity between unrolled networks and distribution-learning-based methods has not been thoroughly investigated; in this study, we propose a novel concept in which we treat naive sampling results from SMLD methods as guidance and unrolled cascades as denoising and refining processes.
3 Background
@@
$\mathcal{E}(x)=(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...\mathcal{S}_{C}x)=(x_{1},x_{2},...,x_{C}),i=1,2,...,C$ (3)
@@
$\mathcal{R}(x_{1},x_{2},...,x_{C})=\sum_{i=1}^{C}\mathcal{S}_{i}^{*}x_{i}$ (4)
@@
It is ill-posed to recover missing data directly from $y$. Hence, a regularization term $\lambda\mathcal{U}(x)$ is routinely added to narrow the solution space, which is
$\hat{x}=\mathop{\arg\min}\limits_{x}\|\mathcal{P}\mathcal{F}\mathcal{E}x-y\|_{2}^{2}+\lambda\mathcal{U}(x)$ (6)
For example, $\lambda\Psi(x)$ is a common regularization term that imposes the sparsity constraint in CS-MRI. Then, gradient descent can be used to solve the problem iteratively as follows:
@@
where $\alpha^{t}$ is the step size in the $t$th iteration and $\psi(x^{t})$ is the gradient of $\Psi(x^{t})$.
4 Methods
Figure 2: (a) The workflow of the proposed SGM-Net. (b) The network structure of the denoising module and the SIE and CIE blocks. (c) The network structure of the $(k+1)$th GIC.
We proposed a score-based generative priors guided model-driven Network (SGM-Net), which is illustrated in Fig. 2. The whole workflow is illustrated in Fig. 2 (a), which comprises three parts: sampling (green components), denoising (blue components), and guided reconstruction (yellow components). The preliminary samples (that is, $x_{T}$ in Fig. 2 (a)) were first obtained using Langevin dynamics and then used as PGIs for the follow-up steps. The gradients of the data distribution were obtained using a PSN. Instead of carefully training and tuning the SMLD process to obtain high-quality samples, we sampled $x_{T}$ from an out-of-distribution pretrained network with fitted hyperparameters. Then, we treated the artifacts in $x_{T}$ as removable and proposed a DM to enhance $x_{T}$. We designed a SIE and a CIE to extract features in DMs, as shown in Fig. 2 (b). Finally, DGIs were used to guide an unrolled network training, as shown in Fig. 2 (c). Details of the proposed method are presented in the following subsections.
4.1 Score matching and Langevin dynamics
The first step of the proposed workflow is to obtain PGIs using SMLD method. In this study, we adopted a PSN proposed in [22], which was successfully extended to MRI reconstruction in [24]. To improve the generality of our workflow, the score network was pretrained with out-of-distribution data, and the tedious tuning and retraining steps were bypassed. The network, remarked as $s_{\theta}$, is a RefineNet [68] with dilated convolutions and conditional instance normalization, which is used to approach the score function of $p(x)$ (that is, $\bigtriangledown_{x}\log p(x)$) and $s_{\theta}\approx\bigtriangledown_{x}\log p(x)$. $s_{\theta}$ is a six-layer U-shaped network. The encoders are designed based on ResNet blocks and the decoders are based on RefineNet blocks.
The annealed Langevin method-based sampling process in MRI can be represented as